OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 13.9 feet b) 12.2 feet c) 7.6 feet d) 8.3 feet

OpenStudy (anonymous):

magic formula for vertex is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

maximum and minimums of parabolas occur at the vertex. x=-b/2a...then plug in for h

OpenStudy (anonymous):

in this case \[-\frac{32}{2\times -16}=1\]